Susan Williams runs a small Flagstaff job shop where garments are made. The job shop employs eight workers. Each worker is paid $12

per hour. During the first week of​ March, each worker worked 45 hours.​ Together, they produced a batch of 156

garments. Of these​ garments, 52 were​ 'seconds' (meaning that they were​ flawed). The seconds were sold for ​$90

each at a factory outlet store. The remaining 104 garments were sold to retail outlets at a price of $210 per garment. What was the labor​ productivity, in dollars per​ labor-hour, at this job shop during the first week of​ March?

Labor productivity​ = __________ dollars/labor hour ​(round your response to two decimal​ places).
Answer #1

Total Number of workers = 8

Number of hours worked by each worker in a week = 45 hours

Total worker hours per week = 45*8 = 360 hours

Wages/hour = $12

Total Wages paid = 360*12 = $4320

Selling Price of 52 garments = $90

Selling price of 104 garments = $210

Total Selling price = 90*52 + 210*104 = $26520

Labor Productivity = Total Selling Price / Total hours per week = 26520 /360 = $73.66 dollars per labor-hour